linear equation by substitution method
a-b=4
b=2-5a
I need the steps to understand not just the answer thanks for your help

Answer by checkley79(3341)

a-b=4
b=2-5a
A-B=4
5A+B=2 ADD
-----------------
6A=6
A=6/6
A=1 ANS.
1-B=4
-B=4-1
-B=3
B=-3 ANS.
PROOF:
-3=2-5*1
-3=2-5
-3=-3